Three liquids are at temperatures of 12 ◦C,
25◦C, and 38◦C, respectively. Equal masses of
the first two liquids are mixed, and the equilibrium temperature is 15◦C. Equal masses of
the second and third are then mixed, and the
the equilibrium temperature is 30.5
◦C.
Find the equilibrium temperature when
equal masses of the first and third are mixed.
Answer in units of ◦C.
1
Expert's answer
2020-04-27T10:15:25-0400

m - mass of each liquids (all masses are equal ).

When the first two are mixed:


"mC_1(T_1\u2212T _{\n12}\n\u200b\t\n )+mC_2(T_2\u2212T _{\n12}\n\u200b\t\n )=0"

"C_1(12\u221215\n\u200b\t\n )+C_2(25\u221215\n\u200b\t\n )=0\\to C_2=0.3C_1"

When the second and therd are mixed:


"mC_3(T_3\u2212T _{\n32}\n\u200b\t\n )+mC_2(T_2\u2212T _{\n32}\n\u200b\t\n )=0"

"C_3(38\u221230.5\n\u200b\t\n )+C_2(25\u221230.5\n\u200b\t\n )=0\\to C_2=1.36C_3"

When the first and therd are mixed:


"mC_1(T_1\u2212T _{\n13}\n\u200b\t\n )+mC_3(T_3\u2212T _{\n13}\n\u200b\t\n )=0"

"\frac{1}{0.3}(15-T_{13})+\frac{1}{1.36}(38-T_{13})=0"

"T_{13}=19.2\degree C"